OpenStudy (anonymous):

what thes volume of 12m by 10 m by 12m cube

Parth (parthkohli):

That doesn't even seem like a cube; a cube has all dimensions equal.

Parth (parthkohli):

Maybe it's a cuboid. In that case, multiply 12 * 10 * 12 to get the volume in m^3.

OpenStudy (zzr0ck3r):

V = L*W*H v = volume l = length h = height w = width

OpenStudy (anonymous):

\[V=l\times w\times h\]\[V=12\times10\times12\]\[V=1440m^3\]

OpenStudy (anonymous):

yes u might have typed wrong but if we take it as a cuboid,then the volume must be l*b*h where l is length,b is breadth and h is height of the cuboid..that is 12*10*12=1440m^3 always remember the unit..always